Mar 12, 2015 Calorimetry

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 11 a.m.

  1. To join via a Web Browser, go to the page [1]
  2. To join via Polycom room system go to the IP Address: ( and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Action Items

  1. BCAL pedestals: Tegan (UofR) will summarize the determination of pedestals in a note with plots in preparation for the collaboration meeting.
  2. Wish-list of action items: discuss attached document (Media:BCAL_Action_Items-2015.pdf) and add chosen items to Wiki page. Zisis will update the list
  3. Will: Create histograms to determine number of pi0s and pi0 width for online monitoring.
  4. Work with Hovanes to activate VME LED pulser setup for FCAL (Elton)


  1. FCAL HDFCAL log book
  2. BCAL HDBCAL log book

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Run conditions
    1. BCAL temperature, overbias settings (Elton)
    2. Data with mode8
    3. HV updates for FCAL (Manuel)
    4. Efficiency in BCal Bias Scan (Foda)[2]
  4. Reconstruction and Simulation
    1. Changes to reconstruction with single-ended hits (Mark)
  5. Calibration
    1. Feb LED runs (Noemi)[3]
    2. pi0 calibration (Will)
    3. Amount of data needed for calibration (Elton)
    4. pi0 and electron calibration (Adesh)[4]
    5. Pedestals (Tegan)
  6. Any other business


Attendees: Elton, Adesh, Will (L207); Manuel (Hall D); Sean (NW); Zisis, Tegan, Noemi, Ahmed (UofR)

  1. Announcements
    • Accelerator close to delivering 3-beam operation
    • Beam loss monitors are currently preventing beam delivery to Hall A at five passes.
    • Solenoid is at 1200 A
    • Hoping for beam this evening.
  2. Action Items
  3. Run conditions
    1. BCAL temperature, overbias settings (Elton)
      • T=10degC for chiller (~11-12 degC on SiPMs), voltage setting for 12 degC, overbias setting is 1.4 V
    2. Data with mode8. Most data will be taken in mode7.
      • Expect to take about 1 hour of mode 8 data per shift
      • Need to check that mode 7 data quality is similar to mode 8 running.
    3. HV updates for FCAL (Manuel)
      • New EPICS GUI displays 100 pmt voltages at a time (one string).
      • Granularity gives better impression of status of FCAL than previous GUI. However, still with several dozen pmts off, most of the FCAL shows red.
      • At the top right the number of pmts not enabled is listed (typically couple of dozen) and updates regularly. This is currently the best diagnostic of the FCAL status.
      • A significant problem occurs now that the FCAL Anagate system is connected to EPICS. The EPICS IOC reboots every couple of hours. This in turn causes a couple of pmts to be turned off. The consequence is that over time, many pmts are found off.
      • Manuel and Hovanes are trying to find the root cause of the reboots.
    4. Efficiency in BCal Bias Scan (Foda)[5]
      • Efficiency was determined for runs with different bias 0.9, 1.2 and 1.4 V. The efficiency increases for nominal clusters by 4-5% for 1.4 V relative to 1.2 V. (Increase for "enhanced efficiency" is only about 1%).
      • No time variation is seen in the (limited) data taken, but temperature was unusually constant during this period.
      • Based on these measurements it was decided to try running at 1.4 V above bias. Threshold scans at 1.2 and 1.4 V should be taken checking efficiency as well as quantity of data.
    5. Other: Zisis inquired about possible improvements in AC unit. Elton: Tim is working with plant services and a) diffusing cold air directed at BCAL crates b) adding temperature sensors at crate c) evaluating optimization of AC operating modes.
  4. Reconstruction and Simulation
    1. Changes to reconstruction with single-ended hits (Mark)
      • Elton: Mark (not present) has investigated what is needed to include single-ended hits into the cluster algorithm and the effort may take time and needs substantial checking.
  5. Calibration
    1. Feb LED runs (Noemi)[6]
      • Updated the group on the work she completed while at JLab. This included getting the LED pulser to work and check that all ADC and TDC channels are now working.
      • The TDC script had to be modified to take a single hit instead of multiple ones. Further updates can check the timing of the pulses and select them that way.
    2. pi0 calibration (Will)
      • Log Entry 3324355
      • New calibration constants have been prepared and will be uploaded to the DB so that they can be used by the upcoming run period.
      • Plots of number of pi0s and signal/background as a function of total energy in the BCAL will be added to log book. These can be useful in making decisions regarding trigger settings.
      • Will is adding a plugin to monitor the pi0 mass peak online. Elton is adding a plugin to monitor the efficiency online.
    3. Amount of data needed for calibration (Elton)
      • GlueX-doc-2677
      • Reviewed the logic behind the amount of data that is needed for calibration of BCAL and FCAL, presented at the RC meeting on Monday. Estimated time is 1-2 weeks for 10x statistics from last run.
    4. pi0 and electron calibration (Adesh)[7]
      • E/p plotted vs p. Select band around E/p~1 as electrons
      • Elton: Why is the band decreasing toward the minimum ionizing level (looks like E/p and E plots are interchanged)? Not clear.
      • Take track projections to FCAL and use sum of 25 blocks (e25) surrounding projected impact to use as estimator of shower energy
      • Use e25/p as a estimator of FCAL gain for blocks within 60 cm of beamline.
      • These calibration constants were used to check uniformity of LED data and width of pi0 peak.
      • It appears the new constants improve the quality slightly relative to uncalibrated.
      • Elton: How does this compare to constants determined by minimizing the pi0 width? Not clear, but will investigate.
      • Sean: Would like the best set of constants available to the loaded into the dB by tomorrow as baseline for the upcoming run. Adesh will use e25/p gains for inner blocks and LED uniformity gains for outer blocks and load db by tomorrow.
    5. Pedestals (Tegan)
      • See Log Entry 3325545 (Log Entry Added after meeting)
      • Determined pedestal widths at 12, 18 and 24 degrees. There is a general trend of an increase in average sigma of between 0.05 and 0.10 in going from 12 degrees to 24 degrees.
  6. Any other business